4x+5y=10 3x-8y=-77.1 solve using addition method or substitution method. I get examples that show them multiplying by 40 and -30 on other problems, but I am not told why they picked those numbers so I am at a loss as where to start.
Answer by stanbon(75887)

4x+5y=10
3x-8y=-77
---------------
It's your choice.
To get rid of the x terms, multiply the
1st equation by 3 and the 2nd equation by -4
-----
To get rid of the y-terms, multiply the 1st equation
by 8 and the 2nd equation by 5.
-----------------------------------
Following the 2nd choice I get:
32x + 40y = 80
15x -40y = -385
--------------------
Add and solve for "x":
47x = -305
---
x = -05/47
----
Substitute and solve for "y":
4x+5y=10
4(-305/47)+ 5y = 10
4(-61/47)+y = 2
4*-61+ 47y = 94
47y = 338
y = 338/47
==================
Cheers,
Stan H.
==================
===================
Cheers,
Stan H.